A) find linear equation to approximate CPI y in year x, with x=0 corresponding to 1993 when CPI is 144.5 in 1993 & 184.2 in 2003 b) Estimate CPI in 2007 C) What rate is CPI increasing?
factor & solve: x2-5x+6=0, 5x2+17x+6=0 & x2-7x-30=0
Thanks.

Answer by ewatrrr(24785)


Hi,
*Note: Note using y = mx + b
x=0 corresponding to 1993 when CPI is 144.5 in 1993
y = mx + 144.5
m = 39.7/10 (change in y over change in x)
y = 3.97x + 144.5
Estimate CPI in 2007 Note: 14 years have passed
y = 3.97*14 + 144.5
What rate is CPI increasing 3.97/yr beginning in 1993
x^2-5x+6=0
factoring:
representing as a PRODUCT,in this case(x-?)(x-?)...according to signs shown
(x-3)(x-2)works as -3x + -2x = -6x and (-3)*(-2) = 6